OpenStudy (anonymous):

Solve the equation and check your solutions. If there is no solution, say no 4x/(x+1 ) = (4x-3)/(x-2) (\4x^2-8x=4x^2-3x+4x-3/) {4x^2-4x^2}-8x=4x-3 {4x^2-4x^2}-3x+4x-3 Can someone please help me please

OpenStudy (tkhunny):

This was okay \(4x^2-8x=4x^2-3x+4x-3\) The next line was not. Your "-3x" seems to have vanished. Why are you carrying [4x^2 -4x^2]? It's just zero. Get rid of it. In the last line, the Equal Sign disappeared. You must be MUCH more careful.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

4x^2-8x=4x^2-3x+4x-3 0-8x=4x-3x =-3x+4x-3

OpenStudy (tkhunny):

Given: \(\dfrac{4x}{x+1} = \dfrac{4x-3}{x-2}\) You first multiplied by the common denominator (x+1)(x-2) to produce \(4x(x-2) = (4x-3)(x+1)\) This is excellent. You then expanded a few things using the distributive property. \(4x^{2} - 8x = 4x^{2} - 3x + 4x - 3\) That was good. Subtracting \(4x^{2}\) from each side gives: \(- 8x = - 3x + 4x - 3\) <== You didn't get this. Your "-3" disappeared. It came back, later, but that's no good. Simplifying the right hand side \(- 8x = x - 3\) <== You have not yet accomplished this result. You ARE close, but just some little oddities in your notation may be confusing you. Don't be afraid to write the ENTIRE equation at each stage. If a friend tells you that you should get good at it SO THAT you can start to skip steps, you may wish to select new friends.

OpenStudy (anonymous):

you will take x from both sides

OpenStudy (anonymous):

Solve for x over the real numbers: (4 x)/(x+1) = (4 x-3)/(x-2) Multiply both sides by a polynomial to clear fractions. Cross multiply: 4 x (x-2) = (x+1) (4 x-3) Write the quadratic polynomial on the left hand side in standard form. Expand out terms of the left hand side: 4 x^2-8 x = (x+1) (4 x-3) Write the quadratic polynomial on the right hand side in standard form. Expand out terms of the right hand side: 4 x^2-8 x = 4 x^2+x-3 Move everything to the left hand side. Subtract 4 x^2+x-3 from both sides: 3-9 x = 0 Factor the left hand side. Factor constant terms from the left hand side: -3 (3 x-1) = 0 Divide both sides by a constant to simplify the equation. Divide both sides by -3: 3 x-1 = 0 Isolate terms with x to the left hand side. Add 1 to both sides: 3 x = 1 Solve for x. Divide both sides by 3: Answer: | | x = 1/3